# Model-driven Stochastic Trace Clustering

### v1

@@ 3.3 Initialization
The EC algorithm requires initial seed selection, which determines the first traces assigned to each cluster. The simplest approach is random selection, but since these initial seeds remain in their respective clusters, their selection can significantly impact clustering quality. To address this, we propose a more sophisticated ++ initialization (Algorithm 4), inspired by how k-means++ clustering handles initialization [22]. This method enhances seed selection by considering pairwise ER distances between traces. This initialization starts again by picking the first seed at random. Then, the pairwise ER distance between this seed and all other variants (traces) is computed. The pairwise ER is calculated by constructing a DFG from just these two traces, and calculating the average ER scores over both traces. The more similar those two traces are, the lower this score (analogous to edit distance). The next seed is selected randomly, weighted by the square of its pairwise ER distance from the current seed. We then continue similarly for the next seeds (if needed), but now calculate the pairwise ER with each seed for each variant (trace). For each variant, we use the minimum distance, i.e., the distance to the closest seed, as weights when selecting the next seed. We continue until we have a number of seeds equal to our required number of clusters.
@@ 4 Experimental setup
  • How uniformly structured are the clusters from a stochastic perspective?
  • How simple and readable are the discovered process models for each cluster?
  • What impact does the clustering method have on (non-stochastic) fitness and precision.
@@ 4.1 Event logs
In order to test our methods on a variety of different process behavior, we used 8 real-life event logs: the helpdesk event log [23], the road traffic fine management (RTFM) event log [24], the BPIC 2013 challenge event logs (the incidents and closed problems) [25], the hospital billing event log [26], the sepsis event log [27], and BPIC 2012 [28] and 2015 event logs [29]. Table 1 provides an overview of the key characteristics of these event logs.
@@ 4.2 Evaluation metrics
Evaluating clustering methods is inherently challenging, as the quality of clusters depends on downstream tasks. To address this, and to stay close to the end-goal of most trace clustering pipelines in process mining, we employ multiple evaluation metrics related to the quality and readability of the discovered process models in each cluster. First of all, we use the Inductive Miner infrequent variant [8], with a (default) noise threshold of 0.2, to discover a Petri net for each cluster. Using the clustered event log and this Petri net we then calculate different (non-stochastic) conformance checking metrics: namely token-based fitness and precision [10], and alignment-based fitness [11] and precision [12]. Next to this, we also calculate the inverse arc-degree to measure the simplicity of the discovered Petri nets [13]. For these we used the implementation in the Python package pm4py [^1] Version 2.7.8.4 [30]. Next, we also discover a DFG (without filters) for each cluster and calculate the entropic relevance as explained before [5] using our simplified version as described above, as we only have fully fitting process models; providing us with a stochastic conformance checking measure. To reiterate: note that this is an average ER score over all traces in the event log, rather than a sum over all traces. To quantify the structural uncertainty of the graph, we also calculate the total graph entropy (GE) directly with only the DFG. Specifically, we compute the Shannon entropy of edge transitions using the probabilities of going from one node (activity) to another.
$GE=\sum_{u\in V}H_{u}=-\sum_{u\in V}\sum_{v\in V,p(u\to v)>0}p(u\to v)\log_{2}p(u\to v)$
Higher entropy indicates a more unpredictable structure, while lower entropy suggests a more deterministic or structured connectivity pattern [31]. Next to GE, we also introduce graph density (GD) as a measure on our DFGs, which measures how close the graph is to being fully connected. Graph density is calculated by dividing the number of edges actually present (non zero) in the DFG, by the maximum number of possible edges:
$GD=\frac{|\text{Edges}|}{|\text{Activity types}|\cdot(|\text{Activity types}|-1)}$
This provides us with another (non-stochastic) simplicity measure, but this time measuring the DFG’s connectivity.
@@
Figure 3: High level overview of the experimental setup.
@@ 4.3 Baseline methods
To assess the performance of our proposed methods, we compare them with several baseline and existing clustering approaches. First, we include two simple baselines. The first baseline does not apply any clustering, meaning that the entire event log is used when calculating the evaluation metrics. The second baseline applies a random clustering approach, where traces are randomly assigned to clusters. Beyond these baselines, we also compare our methods with a frequency-based clustering approach. In this method, we construct a one-hot-encoded vector for each trace, where each dimension of the vector corresponds to an activity type and the values represent the frequency of occurrences of these activities. The vectors are then normalized by dividing each value by the total number of activities in the respective trace. Once the vectors are prepared, clustering is performed using the k-means++ algorithm [22]. We also compare our methods to an embedding-based clustering method using trace2vec [17]. Trace2vec is a self-supervised doc2vec-style approach [32], where trace embeddings are learned by training a shallow neural network to predict the context of activities within a sequence. In our implementation, we use a distributed bag-of-words (PV-DBOW) model trained for 200 epochs. The embedding dimension is set to the square root of the number of unique activity types, while the window size is set to two activities. We use the Gensim library [^2] Version 4.3.2 [33]. After obtaining the trace embeddings, we again apply the k-means++ algorithm to obtain clusters. We further compare our approach with the active trace clustering method ActiTraC, in both its frequency-based and distance-based variant. For both ActiTraC methods we use the implementation within the ProM framework [^3] Version 6.13 [34], and the default parameter settings, where the minimal cluster size is set to 0.25, the target ICS-fitness is set to 1.0, and the heuristic net discovery is used. To maintain consistency across all methods, the number of clusters is set to be the same as in the other approaches. Any traces that are not initially assigned to clusters are distributed among the existing clusters.
@@ 4.4 Elbow experiment
Since entropic clustering requires a predefined number of clusters, we conduct an elbow experiment to determine a reasonable cluster count. In this experiment, we plot the values of entropic relevance, graph density, and graph entropy across different numbers of clusters. We perform this analysis for the efficient random clustering, frequency-based clustering, EC, and ECsplit (both with ++ initialization). The results of these experiments can be found in the online repository. For some event logs, a clear elbow point with consensus among metrics and methods could be achieved, but for others this was more difficult. Nevertheless, since our downstream experiments focus on comparing different clustering methods under the same fixed number of clusters, a potentially arbitrary choice of cluster count does not significantly impact the experimental validity. The specific number of clusters selected for each event log is provided in Table 1.
@@ 4.5 Ranking
To evaluate the effectiveness of each clustering method, we determine its ranking for each event log along each metric. For each metric, we then calculate the average rank of each clustering method. The statistical significance of the ranking is determined following the well-established methodology described by Demšar et al. [35]. Specifically, for each metric, we perform Friedman’s test to assess whether the ranking of clustering methods is statistically significant across datasets. Friedman’s test is a non-parametric test that uses the average ranks of clustering methods across datasets. The null hypothesis assumes that all methods perform equivalently on average (i.e., their ranks should be equal). The Friedman test statistic follows a chi-square distribution with $\chi^{2}_{F}$ and $k-1$ degrees of freedom, where $k$ is the number of clustering methods. If $p<0.05$, we reject the null hypothesis, indicating that at least one method differs significantly. If Friedman’s test is significant, we proceed with post-hoc pairwise comparisons using the Nemenyi test, which determines whether the performance differences between clustering methods are statistically meaningful, and accounts for multiple comparisons. The results from the Nemenyi test are further visualized using a Critical Difference (CD) diagram, which provides an intuitive way to interpret the findings. In a CD diagram, clustering methods are placed along a ranking scale, and those that are not significantly different are connected with a horizontal line.
@@ 5 Experimental results
For each event log, the same number of clusters is extracted using each clustering method. The different metrics, as described above, were computed for each cluster, alongside the metric values for the full event log, which served as a baseline. The final score for each method is determined by taking the weighted average of the scores, where the weight is based on the number of cases in each cluster, ensuring that each case is weighted equally. The detailed results for each event log can be found in Appendix A. Due to the experiments running out of time the alignments-based conformance checking methods were not used for logs BPIC12 and BPIC15. The full results, for each separate cluster, can be found online. Table 2 presents the average rankings of each clustering method across the eight event logs, evaluated separately for each metric. The table uses a color-coded ranking system, where green represents the highest rank (1) and white represents the lowest (12). The best-performing method for each metric is highlighted in bold with an underline. The Friedman $\chi^{2}$ statistic and corresponding p-value indicate whether at least one method significantly outperforms the others for a given metric. Statistically significant p-values (i.e., below 0.05) are marked in bold, which in this case applies only to simplicity, ER, graph density, and entropy. Notably, the proposed EC methods outperform the rest on these metrics.
@@ 6 Discussion & limitations
In our current experimental setup, we did not include other stochastic conformance checking metrics. Other conformance checking metrics, such as using the Earth mover’s distance [14], would be most meaningful if we were not using fully fitting (non-filtered) DFGs as the discovered model. We would therefore have to decide on one stochastic discovery technique, which could have a large influence on the end results. Or, to properly account for this, try different techniques (and filter) leading to exponentially more experiments. We opted for ER as our primary measure because it captures the (stochastic) the lack of structure in clusters in a general way, independent of the possible downstream discovery.
@@
Unlike methods such as k-means, our approach does not rely on shifting cluster centroids to reach an equilibrium, as defining efficient cluster centers or medoids for process behavior is not straightforward. Consequently, the EC method could still remain somewhat sensitive to initialization in its current form. Our ++ and ++norm initializations still select the first seed randomly, which can lead to suboptimal starting conditions, and for the recursive ECsplit variant at each split. A more computationally expensive approach, incorporating all pairwise distances, could be explored to mitigate this issue. Additionally, the ++ and ++norm initializations, despite the latter’s normalization to account for loops, tend to favor outlier traces as initial seeds. While this may be beneficial in some cases, it could be suboptimal for certain applications. The lack of statistically significant differences between initialization methods may point to a lower-than-expected sensitivity to initialization. One might consider using the sum of all trace ER scores per cluster, rather than the average, to better reflect the the actual information compression cost. However, this choice does not affect any of the conclusions of this work. For our clustering algorithm, the impact would be negligible, as it relies solely on the ER scores of candidate traces. While the evaluation scores might differ slightly due to this change, we employ weighted averages, ensuring that the relative ranking of the algorithms remains unchanged.
@@ 7 Conclusion & future work
For future work, a larger-scale experiment could be conducted, potentially using controlled synthetic data. This would allow us to systematically analyze which types of process behaviors and challenges affect different clustering techniques. A simple but useful extension could involve introducing a robust version of the algorithm that detects and handles outliers, as they likely skew clustering results in the current approach. Further extensions could expand the model-driven stochastic clustering paradigm by incorporating other process model types, discovery algorithms, and stochastic conformance checking techniques. We opted for the current combination due to its efficiency and widespread adoption in commercial tools. Our algorithm (see Alg. 2) updates each cluster’s DFG incrementally when assigning traces. An alternative approach could forgo this update and use the DFGs as they are, relying on an ER metric implementation [5] including an appropriate background model for non-fitting traces. Another avenue for improvement lies in the initialization phase. Seed selection could be guided by expert knowledge, if available. Given the scalability of our method, running the clustering multiple times and selecting the best result could help mitigate volatility due to initialization. Additionally, since determining the correct number of clusters upfront is often impractical, a variant that dynamically adjusts the number of clusters, based on a maximum ER threshold, for example, could be explored. Finally, our approach could be extended to incorporate additional process features, such as the resource performing an activity or sensor measurements. One way to achieve this would be to adjust probability calculations based on these features. For categorical attributes, probabilities could be derived from frequency distributions in the current traces, while for continuous attributes, they could be estimated using a fitted (normal) distribution.

### v2

@@ 4.3 Initialization
The EC algorithm requires initial seed selection, which determines the first traces assigned to each cluster. The simplest approach is random selection, but since these initial seeds remain in their respective clusters, their selection can significantly impact clustering quality. To address this, we propose a more sophisticated ++ initialization (Algorithm 4), inspired by how k-means++ clustering handles initialization [Arthur2007]. This method enhances seed selection by considering pairwise ER distances between traces. This initialization starts again by picking the first seed at random. Then, the pairwise ER distance between this seed and all other variants (traces) is computed. The pairwise ER is calculated by constructing a DFG from just these two traces, and calculating the average ER scores over both traces. The more similar those two traces are, the lower this score (analogous to edit distance). The next seed is selected randomly, weighted by the square of its pairwise ER distance from the current seed. We then continue similarly for the next seeds (if needed), but now calculate the pairwise ER with each seed for each variant (trace). For each variant, we use the minimum distance, i.e., the distance to the closest seed, as weights when selecting the next seed. We continue until we have a number of seeds equal to our required number of clusters.
@@ 5 Experimental setup
  • RQ1: How uniformly structured are the clusters from a stochastic perspective?
  • RQ2: How simple and readable are the discovered process models for each cluster?
  • RQ3: What impact does the clustering method have on (non-stochastic) fitness and precision.
@@ 5.1 Event logs
In order to test our methods on a variety of different process behavior, we used 8 real-life event logs: the helpdesk event log [Helpdesk], the road traffic fine management (RTFM) event log [RTFM], the BPIC 2013 challenge event logs (the incidents and closed problems) [BPIC13], the hospital billing event log [hospitalbilling], the sepsis event log [Sepsis], and BPIC 2012 [BPIC12] and 2015 event logs [BPIC15]. Table 1 provides an overview of the key characteristics of these event logs. Note that for BPIC15 log we use a version where all 5 municipality logs are merged into one, as done before in other works concerning clustering [DeKoninck2018].
@@ 5.2 Evaluation metrics
Evaluating clustering methods is inherently challenging because the quality of clusters depends on the downstream task. To address this, and to stay close to the end-goal of most trace clustering pipelines in process mining, we employ multiple evaluation metrics related to the quality and readability of the discovered process models in each cluster. To align our evaluation with the typical end-goal of trace clustering in process mining, we use a diverse set of metrics that jointly assess three aspects: stochastic structure, model simplicity, and non-stochastic conformance. These aspects correspond to our research questions RQ1-RQ3 and are reflected in the metrics described below.
First, to evaluate non-stochastic conformance and model quality within each cluster, we use the Inductive Miner infrequent variant [Leemans2014] with its default noise threshold of 0.2 to discover a Petri net for each cluster. Using the clustered event log and the corresponding Petri net, we then compute several different (non-stochastic) conformance checking metrics: namely, token-based fitness and precision [rozinat2008], and alignment-based fitness [Vanderaalst2012] and precision [Adriansyah2015]. These metrics quantify how well the discovered model represents the behavior in the cluster and how much the clustering itself can improve these conformance measures when using the same discovery technique, which directly addresses RQ3. To capture model simplicity, we calculate the inverse arc-degree of the discovered Petri nets  [SanchezGonzalez2010], which serves as a structural simplicity measure and addresses RQ2. For these metrics, we used the implementation available in the Python package pm4py (Version 2.7.8.4)  [Berti2023].
Second, to evaluate how uniformly structured the clusters are from a stochastic perspective (RQ1), we discover an unfiltered directly-follows graph (DFG) for each cluster and compute the entropic relevance (ER) [Alkhammash2022], following Algorithm 1. We use a simplified ER variant that is appropriate in our setting because the discovered process models are fully fitting. ER acts as a stochastic conformance measure: high ER values indicate greater structural homogeneity within the cluster. We report both the total ER of the log-DFG combination (ERsum), as proposed in [Alkhammash2022], and the average ER score over all traces in the event log (ERav).
To quantify the structural uncertainty of the DFG, we also calculate the total graph entropy (graphentr) directly on the DFG, again addressing RQ1. Specifically, we calculate the Shannon entropy of edge transitions based on the probabilities of moving from one node (activity) to another. Higher entropy indicates a more unpredictable structure, while lower entropy suggests more deterministic and structured connectivity patterns [Dehmer2011].
$\texttt{graph\textsubscript{entr}}=\sum_{u\in V}H_{u}=-\sum_{u\in V}\sum_{v\in V,p(u\to v)>0}p(u\to v)\log_{2}p(u\to v)$
Finally, to assess an additional aspect of model simplicity and readability (RQ2), we calculate the graph density (graphdens) of each DFG. Graph density is obtained by dividing the number of observed edges in the DFG by the maximum possible number of edges between activity types:
$\texttt{graph\textsubscript{dens}}=\frac{|\text{Edges}|}{|\text{Activity types}|\cdot(|\text{Activity types}|-1)}$
Graphdens quantifies how close the DFG is to being fully connected and provides a complementary non-stochastic simplicity measure that focuses on the DFG’s connectivity patterns.
@@
Figure 3: High-level overview of the experimental setup.
@@ 5.3 Baseline methods
To assess the performance of our proposed methods, we compare them with several baseline and existing clustering approaches. First, we include two simple baselines. The first baseline does not apply any clustering, meaning that the entire event log is used when calculating the evaluation metrics. The second baseline applies a random clustering approach, where traces are randomly assigned to clusters. Beyond these baselines, we also compare our methods with a frequency-based clustering approach. In this method, we construct a frequency encoded vector for each trace, where each dimension of the vector corresponds to an activity type and the values represent the frequency of occurrences of these activities. The vectors are then normalized by dividing each value by the total number of activities in the respective trace. Once the vectors are prepared, clustering is performed using the k-means++ algorithm [Arthur2007]. We also compare our methods to an embedding-based clustering method using trace2vec [DeKoninck2018]. Trace2vec is a self-supervised doc2vec-style approach [Le2014], where trace embeddings are learned by training a shallow neural network to predict the context of activities within a sequence. In our implementation, we use a distributed bag-of-words (PV-DBOW) model trained for 200 epochs. The embedding dimension is set to the square root of the number of unique activity types, while the window size is set to two activities. We use the Gensim library (Version 4.3.2) [gensim]. After obtaining the trace embeddings, we again apply the k-means++ algorithm to obtain clusters. We further compare our approach with the active trace clustering method ActiTraC, in both its frequency-based and distance-based variant. For both ActiTraC methods we use the implementation within the ProM framework (Version 6.13) [ProM], and the default parameter settings, where the minimal cluster size is set to 0.25, the target ICS-fitness is set to 1.0, and the heuristic net discovery is used. To maintain consistency across all methods, the number of clusters is set to be the same as in the other approaches. Any traces that are not initially assigned to clusters are distributed among the existing clusters.
@@ 5.4 Elbow experiment
Since entropic clustering (like many other trace clustering methods) requires a predefined number of clusters, we first performed an independent analysis to determine a suitable cluster count $k$ for each event log. In this preliminary parameter determination phase, we conducted an “elbow experiment” by plotting the values of entropic relevance, graph density, and graph entropy across a range of cluster counts (from $k=1$ to $k=10$). These metrics were selected based on their computational efficiency. We analyzed the trends for random clustering, frequency-based clustering, EC, and ECsplit (both with ++ initialization), also selected on their computational efficiency.
The detailed plots from this analysis are presented in B. While some event logs exhibited a distinct elbow point with consensus across metrics, others required a more qualitative selection. Crucially, this selection was made prior to the comparative evaluation. Once a specific $k$ was selected for an event log (as listed in Table 1), it was fixed and applied to all clustering methods in the subsequent experiments. This ensures that our downstream evaluation compares different clustering methods on a level playing field: any potential arbitrariness in the choice of $k$ affects all methods equally and thus does not bias the relative ranking of the algorithms.
@@ 5.5 Ranking
To evaluate the effectiveness of each clustering method, we determine its ranking for each event log along each metric. For each metric, we then calculate the average rank of each clustering method. The statistical significance of the ranking is determined following the well-established methodology described by Demšar et al. [Demsar2006]. Specifically, for each metric, we perform Friedman’s test to assess whether the ranking of clustering methods is statistically significant across datasets. Friedman’s test is a non-parametric test that uses the average ranks of clustering methods across datasets. The null hypothesis assumes that all methods perform equivalently on average (i.e., their ranks should be equal). The Friedman test statistic follows a chi-square distribution with $\chi^{2}_{F}$ and k1k-1 degrees of freedom, where $k$ is the number of clustering methods. If $p<0.05$, we reject the null hypothesis, indicating that at least one method differs significantly.
Because we run a separate Friedman test for each evaluation metric (multiple tests), we control the family-wise Type I error across metrics by applying the Holm step-down correction to the Friedman $p$-values. [Holm1979]. Holm’s procedure works by sorting the $p$-values and multiplying each by $(m-\text{rank}+1)$ where $m$ indicates the total number of metrics (8 in this case). It also enforces monotonicity: adjusted $p$-values must be non-decreasing, meaning that each corrected value is at least as large as those for all lower-ranked tests. This correction is somewhat conservative: one could reasonably argue that the metrics do not form a single fully coherent family, as some are related while others (e.g., fitness versus precision) measure very different aspects. Moreover, such corrections are most relevant when a claim hinges on any individual null hypothesis being rejected, i.e., if you treat each metric as a qualitatively distinct evaluation dimension. In our case, we are more interested in how methods compare on each metric individually than in whether a method is superior on at least one metric. Nevertheless, since some of our metrics do correlate to a certain extent, we apply Holm’s correction for completeness.
If the Holm-adjusted p-value is below 0.05, we reject the null hypothesis for that metric, indicating that at least one method differs significantly. We then proceed with post-hoc pairwise comparisons using the Nemenyi test, which determines whether the performance differences between clustering methods are statistically meaningful, and accounts for multiple comparisons. The results from the Nemenyi test are further visualized using a Critical Difference (CD) diagram, which provides an intuitive way to interpret the findings. In a CD diagram, clustering methods are placed along a ranking scale, and those that are not significantly different are connected with a horizontal line.
@@ 6 Experimental results
For each event log, the same number of clusters is extracted using each clustering method. The different metrics, as described above, were computed for each cluster, alongside the metric values for the full event log, which served as a baseline. The final score for each method is determined by taking the weighted average of the scores, where the weight is based on the number of cases in each cluster, ensuring that each case is weighted equally. The detailed results for each event log can be found in  A. Due to the experiments running out of time, the alignments-based conformance checking methods were not used for logs BPIC12 and BPIC15. The full results, for each separate cluster, can be found online. Table 2 presents the average rankings of each clustering method across the eight event logs, evaluated separately for each metric. The table uses a color-coded ranking system, where green represents the highest rank (1) and white represents the lowest (12). Since the rankings for ERav and ERsum are the same, we only add ER once. The best-performing method for each metric is highlighted in bold with an underline. The Friedman $\chi^{2}$ statistic and both the raw and Holm-adjusted $p$-values indicate whether at least one method significantly outperforms the others for a given metric. Statistically significant p-values (i.e., below 0.05) are marked in bold, which in this case applies only to simplicity, ER, graph density, and graph entropy. Notably, the proposed EC methods outperform the rest on these metrics.
@@ 7 Discussion & limitations
Based on these findings, we can offer practical guidance on when to apply Entropic Clustering versus traditional approaches. If the primary analytical goal is to produce precise, replayable Petri nets where structural constraints (such as parallelism and exact traces) are paramount, traditional frequency-based or vector-based clustering methods may be more suitable. However, Entropic Clustering is the superior choice when the analysis relies on Directly-Follows Graphs, as is common in commercial process mining tools, and when the user requires an understanding of the dominant probabilistic behaviors. By penalizing low-probability transitions, EC is particularly effective at untangling “spaghetti” processes into coherent, stochastically valid subgroups, making it ideal for exploratory analysis and interpretability-focused tasks.
In our current experimental setup, we did not include other stochastic conformance checking metrics. Other conformance checking metrics, such as using the Earth mover’s distance [Leemans2019], would be most meaningful if we were not using fully fitting (non-filtered) DFGs as the discovered model. We would therefore have to decide on one stochastic discovery technique, which could have a large influence on the end results. Or, to properly account for this, try different techniques (and filter) leading to exponentially more experiments. We opted for ER as our primary measure because it captures the (stochastic) the lack of structure in clusters in a general way, independent of the possible downstream discovery.
Regarding the computation of this metric, we recognize that a lower numerical cut-off (allowing smaller probabilities) would produce more precise absolute ER scores for some logs. However, our experiments are designed to compare the relative performance of clustering methods. Since this probability floor $\varepsilon$ is applied uniformly across all methods, it acts as a consistent penalty for extreme outliers and is unlikely to change the relative ranking of the algorithms or the overall conclusions.
@@
Unlike methods such as k-means, our approach does not rely on shifting cluster centroids to reach an equilibrium, as defining efficient cluster centers or medoids for process behavior is not straightforward. Consequently, the EC method could still remain somewhat sensitive to initialization in its current form. Our ++ and ++norm initializations still select the first seed randomly, which can lead to suboptimal starting conditions, and for the recursive ECsplit variant at each split. A more computationally expensive approach, incorporating all pairwise distances, could be explored to mitigate this issue. Additionally, the ++ and ++norm initializations, despite the latter’s normalization to account for loops, tend to favor outlier traces as initial seeds. While this may be beneficial in some cases, it could be suboptimal for certain applications. The lack of statistically significant differences between initialization methods may point to a lower-than-expected sensitivity to initialization. Nevertheless, it might be interesting to experiment with a k-medoids-style approach, where the initialisation is updated over multiple runs. This could be done by, e.g., looking at the trace with the highest play-out probability (or lowest ER) with the current cluster’s DFG, in each cluster.
@@ 8 Conclusion & future work
For future work, a larger-scale experiment could be conducted, potentially using controlled synthetic data. This would allow us to systematically analyze which types of process behaviors and challenges affect different clustering techniques. A simple but useful extension could involve introducing a robust version of the algorithm that detects and handles outliers, as they likely skew clustering results in the current approach. Further extensions could expand the model-driven stochastic clustering paradigm by incorporating other process model types, discovery algorithms, and stochastic conformance checking techniques. We opted for the current combination due to its efficiency and widespread adoption in commercial tools. Our algorithm (see Alg. 2) updates each cluster’s DFG incrementally when assigning traces. An alternative approach could forgo this update and use the DFGs as they are, relying on an ER metric implementation [Alkhammash2022] including an appropriate background model for non-fitting traces. Another avenue for improvement lies in the initialization phase. Seed selection could be guided by expert knowledge, if available. Given the scalability of our method, running the clustering multiple times and selecting the best result could help mitigate volatility due to initialization. This could also be done in a k-medoids-style approach, by running the clustering algorithm multiple times (with updated initialization) until no change is detected. The new initial clusters could in each iteration be selected as the trace with the lowest ER score (with the cluster’s DFG). Additionally, since determining the correct number of clusters upfront is often impractical, a variant that dynamically adjusts the number of clusters, based on a maximum ER threshold, for example, could be explored. Finally, our approach could be extended to incorporate additional process features, such as the resource performing an activity or sensor measurements. One way to achieve this would be to adjust probability calculations based on these features. For categorical attributes, probabilities could be derived from frequency distributions in the current traces, while for continuous attributes, they could be estimated using a fitted (normal) distribution.